# Jack of All Trades, Master of Some, a Multi-Purpose Transformer Agent

### v1

@@ Appendix A Full results
Boss Level 0.06 ± 0.21 0.94 ± 0.05 0.48 ± 0.45 0.48 ± 0.51
Boss Level No Unlock 0.06 ± 0.19 0.94 ± 0.05 0.44 ± 0.45 0.43 ± 0.51
Find Obj S5 0.08 ± 0.23 0.95 ± 0.04 0.95 ± 0.03 1.00 ± 0.04
Go To 0.13 ± 0.29 0.92 ± 0.07 0.80 ± 0.27 0.85 ± 0.35
Go To Door 0.45 ± 0.38 0.99 ± 0.00 0.99 ± 0.01 1.00 ± 0.01
Go To Imp Unlock 0.07 ± 0.22 0.83 ± 0.13 0.50 ± 0.44 0.56 ± 0.59
Go To Local 0.16 ± 0.30 0.93 ± 0.04 0.88 ± 0.14 0.94 ± 0.18
Go To Obj 0.13 ± 0.27 0.93 ± 0.03 0.93 ± 0.04 0.99 ± 0.05
Go To Obj Door 0.53 ± 0.39 0.99 ± 0.01 0.98 ± 0.04 0.97 ± 0.08
Go To Red Ball 0.17 ± 0.30 0.93 ± 0.04 0.91 ± 0.08 0.98 ± 0.11
Go To Red Ball Grey 0.12 ± 0.27 0.92 ± 0.05 0.91 ± 0.06 0.99 ± 0.08
@@
Go To Red Blue Ball 0.12 ± 0.27 0.92 ± 0.05 0.88 ± 0.11 0.96 ± 0.13
Go To Seq 0.08 ± 0.23 0.94 ± 0.05 0.73 ± 0.34 0.75 ± 0.40
Key Corridor 0.00 ± 0.00 0.91 ± 0.01 0.88 ± 0.10 0.97 ± 0.11
Mini Boss Level 0.07 ± 0.21 0.89 ± 0.10 0.69 ± 0.35 0.76 ± 0.43
Move Two Across S8N9 0.00 ± 0.00 0.96 ± 0.01 0.03 ± 0.15 0.03 ± 0.16
One Room S8 0.08 ± 0.21 0.92 ± 0.03 0.92 ± 0.03 1.00 ± 0.04
Open 0.10 ± 0.24 0.95 ± 0.05 0.93 ± 0.11 0.97 ± 0.13
@@
Open Doors Order N4 0.16 ± 0.30 0.99 ± 0.01 0.96 ± 0.11 0.97 ± 0.13
Open Red Door 0.08 ± 0.21 0.92 ± 0.03 0.92 ± 0.02 1.00 ± 0.03
@@
Pickup 0.08 ± 0.22 0.92 ± 0.07 0.72 ± 0.34 0.77 ± 0.40
Pickup Above 0.02 ± 0.09 0.91 ± 0.07 0.92 ± 0.06 1.01 ± 0.07
Pickup Dist 0.10 ± 0.24 0.86 ± 0.21 0.88 ± 0.13 1.03 ± 0.18
Pickup Loc 0.08 ± 0.23 0.91 ± 0.04 0.84 ± 0.20 0.91 ± 0.24
Put Next S7N4 0.00 ± 0.03 0.96 ± 0.01 0.82 ± 0.26 0.86 ± 0.27
Put Next Local 0.00 ± 0.05 0.92 ± 0.03 0.60 ± 0.36 0.65 ± 0.39
Synth 0.11 ± 0.26 0.93 ± 0.06 0.68 ± 0.39 0.69 ± 0.47
Synth Loc 0.13 ± 0.29 0.94 ± 0.06 0.82 ± 0.31 0.85 ± 0.38
Synth Seq 0.07 ± 0.20 0.95 ± 0.04 0.57 ± 0.44 0.57 ± 0.50
Unblock Pickup 0.08 ± 0.22 0.91 ± 0.08 0.76 ± 0.33 0.82 ± 0.39
Unlock 0.03 ± 0.15 0.87 ± 0.10 0.55 ± 0.42 0.63 ± 0.50
@@
Unlock Pickup 0.00 ± 0.00 0.75 ± 0.04 0.76 ± 0.03 1.01 ± 0.04
Unlock To Unlock 0.00 ± 0.00 0.96 ± 0.00 0.86 ± 0.29 0.89 ± 0.30
@@
BinPicking 1.9 ± 0.4 425.6 ± 101.9 374.2 ± 168.2 0.88 ± 0.40
Box Close 76.4 ± 17.9 512.5 ± 107.8 510.1 ± 117.5 0.99 ± 0.27
Button Press 31.7 ± 5.2 643.1 ± 12.8 556.8 ± 198.9 0.86 ± 0.33
Button Press Topdown 29.0 ± 10.4 490.2 ± 27.2 265.2 ± 77.9 0.51 ± 0.17
Button Press Topdown Wall 29.0 ± 10.5 497.2 ± 31.4 260.1 ± 67.7 0.49 ± 0.14
Button Press Wall 9.0 ± 4.0 675.4 ± 15.0 621.7 ± 137.1 0.92 ± 0.21
Coffee Button 31.7 ± 6.4 731.1 ± 29.3 250.5 ± 266.9 0.31 ± 0.38
Coffee Pull 4.1 ± 0.4 259.9 ± 88.5 55.1 ± 97.0 0.20 ± 0.38
Coffee Push 4.2 ± 0.8 496.8 ± 118.2 269.2 ± 237.8 0.54 ± 0.48
Dial Turn 29.6 ± 16.7 793.6 ± 80.1 738.2 ± 168.4 0.93 ± 0.22
Disassemble 40.3 ± 7.5 42.8 ± 6.3 39.1 ± 11.9 -0.47 ± 4.70
Door Close 5.3 ± 1.3 529.7 ± 27.2 528.2 ± 29.9 1.00 ± 0.06
Door Lock 112.3 ± 28.6 811.5 ± 34.1 676.5 ± 192.7 0.81 ± 0.28
Door Open 56.4 ± 11.2 581.9 ± 19.7 572.8 ± 57.5 0.98 ± 0.11
Door Unlock 94.2 ± 15.6 802.9 ± 17.1 654.9 ± 260.6 0.79 ± 0.37
Drawer Close 116.7 ± 253.1 867.9 ± 4.5 663.0 ± 214.5 0.73 ± 0.29
Drawer Open 126.8 ± 25.2 493.0 ± 2.5 489.1 ± 21.3 0.99 ± 0.06
Faucet Close 253.1 ± 22.9 753.9 ± 13.4 361.3 ± 72.3 0.22 ± 0.14
Faucet Open 244.1 ± 23.3 705.8 ± 7.1 637.9 ± 134.5 0.85 ± 0.29
Hammer 95.3 ± 9.0 693.2 ± 34.6 691.7 ± 25.3 1.00 ± 0.04
Hand Insert 2.8 ± 3.5 740.5 ± 36.7 719.6 ± 99.3 0.97 ± 0.13
Handle Press 80.4 ± 110.2 855.9 ± 72.7 731.9 ± 261.9 0.84 ± 0.34
Handle Press Side 57.0 ± 39.5 861.1 ± 20.0 84.3 ± 113.3 0.03 ± 0.14
Handle Pull 10.3 ± 13.5 669.4 ± 24.8 501.3 ± 209.5 0.74 ± 0.32
Handle Pull Side 2.1 ± 2.8 384.7 ± 102.9 233.1 ± 199.5 0.60 ± 0.52
Lever Pull 60.3 ± 15.8 612.0 ± 38.9 250.2 ± 228.6 0.34 ± 0.41
Peg Insert Side 1.7 ± 0.4 315.2 ± 140.1 288.0 ± 157.9 0.91 ± 0.50
Peg Unplug Side 4.7 ± 2.8 456.1 ± 81.7 68.5 ± 125.3 0.14 ± 0.28
@@
Pick Place 1.6 ± 1.0 419.1 ± 98.2 264.2 ± 195.7 0.63 ± 0.47
Pick Place Wall 0.0 ± 0.0 450.6 ± 64.1 6.9 ± 45.0 0.02 ± 0.10
Plate Slide 74.6 ± 13.8 527.0 ± 155.3 497.4 ± 168.7 0.93 ± 0.37
Plate Slide Back 33.5 ± 11.2 718.2 ± 87.4 196.8 ± 1.7 0.24 ± 0.00
Plate Slide Back Side 34.3 ± 11.5 729.6 ± 69.1 697.5 ± 137.8 0.95 ± 0.20
Plate Slide Side 22.6 ± 17.4 662.8 ± 102.8 122.6 ± 24.5 0.16 ± 0.04
Push 5.5 ± 2.4 750.6 ± 44.0 604.2 ± 261.9 0.80 ± 0.35
Push Back 1.2 ± 0.2 85.0 ± 107.1 91.4 ± 115.0 1.08 ± 1.37
Push Wall 6.1 ± 3.2 748.9 ± 10.6 116.5 ± 208.1 0.15 ± 0.28
Reach 149.7 ± 44.7 681.4 ± 133.7 325.3 ± 159.2 0.33 ± 0.30
Reach Wall 143.3 ± 36.6 746.1 ± 104.2 634.6 ± 231.4 0.81 ± 0.38
Shelf Place 0.0 ± 0.0 241.3 ± 24.6 124.6 ± 112.8 0.52 ± 0.47
Soccer 5.7 ± 4.6 375.2 ± 140.2 364.5 ± 175.4 0.97 ± 0.47
Stick Pull 2.6 ± 1.4 523.6 ± 18.9 398.6 ± 205.6 0.76 ± 0.39
Stick Push 2.8 ± 1.0 627.9 ± 10.2 158.3 ± 264.6 0.25 ± 0.42
Sweep 11.2 ± 7.3 494.8 ± 43.3 15.6 ± 9.3 0.01 ± 0.02
Sweep Into 12.5 ± 10.7 799.2 ± 19.1 775.3 ± 119.0 0.97 ± 0.15
Window Close 57.5 ± 7.1 591.3 ± 38.6 423.3 ± 203.9 0.69 ± 0.38
Window Open 43.4 ± 2.1 590.8 ± 57.1 593.1 ± 54.8 1.00 ± 0.10
@@
Ant -59.9 ± 99.6 5846.4 ± 942.6 5268.0 ± 1495.4 0.90 ± 0.25
Inverted Double Pendulum 57.5 ± 17.5 9338.7 ± 352.6 4750.1 ± 931.2 0.51 ± 0.10
Half Cheetah -285.0 ± 79.8 7437.8 ± 173.3 6659.7 ± 409.7 0.90 ± 0.05
Hopper 18.4 ± 17.1 1858.7 ± 534.1 1835.9 ± 532.2 0.99 ± 0.29
Humanoid 122.0 ± 35.3 6281.0 ± 1795.8 697.4 ± 108.1 0.09 ± 0.02
Inverted Pendulum 6.1 ± 3.5 475.4 ± 179.0 116.3 ± 20.2 0.23 ± 0.04
Pusher -149.7 ± 7.4 -25.2 ± 6.7 -26.3 ± 6.3 0.99 ± 0.05
Reacher -43.0 ± 3.9 -5.7 ± 2.5 -6.1 ± 2.6 0.99 ± 0.07
Humanoid Standup 33135.8 ± 2481.9 273574.2 ± 85253.3 118125.2 ± 24880.3 0.35 ± 0.10
Swimmer 0.8 ± 10.7 92.2 ± 4.4 93.3 ± 3.8 1.01 ± 0.04
Walker 2d 2.7 ± 6.1 4631.2 ± 1059.0 4662.4 ± 762.7 1.01 ± 0.16

### v2

@@ Appendix A Full results
Boss Level 0.06 ± 0.21 0.94 ± 0.05 0.52 ± 0.43 0.53 ± 0.49
Boss Level No Unlock 0.06 ± 0.19 0.94 ± 0.05 0.48 ± 0.43 0.48 ± 0.48
Find Obj S5 0.08 ± 0.23 0.95 ± 0.04 0.95 ± 0.04 1.01 ± 0.05
Go To 0.13 ± 0.29 0.92 ± 0.07 0.83 ± 0.27 0.89 ± 0.34
Go To Door 0.45 ± 0.38 0.99 ± 0.00 0.99 ± 0.02 0.99 ± 0.04
Go To Imp Unlock 0.07 ± 0.22 0.83 ± 0.13 0.60 ± 0.41 0.70 ± 0.54
Go To Local 0.16 ± 0.30 0.93 ± 0.04 0.88 ± 0.14 0.94 ± 0.19
Go To Obj 0.13 ± 0.27 0.93 ± 0.03 0.93 ± 0.03 1.00 ± 0.04
Go To Obj Door 0.53 ± 0.39 0.99 ± 0.01 0.96 ± 0.10 0.94 ± 0.21
Go To Red Ball 0.17 ± 0.30 0.93 ± 0.04 0.92 ± 0.05 0.99 ± 0.07
Go To Red Ball Grey 0.12 ± 0.27 0.92 ± 0.05 0.91 ± 0.07 0.99 ± 0.08
@@
Go To Red Blue Ball 0.12 ± 0.27 0.92 ± 0.05 0.88 ± 0.11 0.95 ± 0.14
Go To Seq 0.08 ± 0.23 0.94 ± 0.05 0.72 ± 0.34 0.74 ± 0.40
Key Corridor 0.00 ± 0.00 0.91 ± 0.01 0.86 ± 0.16 0.94 ± 0.18
Mini Boss Level 0.07 ± 0.21 0.89 ± 0.10 0.61 ± 0.39 0.65 ± 0.47
Move Two Across S8N9 0.00 ± 0.00 0.96 ± 0.01 0.02 ± 0.13 0.02 ± 0.13
One Room S8 0.08 ± 0.21 0.92 ± 0.03 0.92 ± 0.04 1.00 ± 0.04
Open 0.10 ± 0.24 0.95 ± 0.05 0.94 ± 0.11 0.98 ± 0.13
@@
Open Doors Order N4 0.16 ± 0.30 0.99 ± 0.01 0.95 ± 0.17 0.95 ± 0.20
Open Red Door 0.08 ± 0.21 0.92 ± 0.03 0.91 ± 0.03 1.00 ± 0.04
@@
Pickup 0.08 ± 0.22 0.92 ± 0.07 0.76 ± 0.32 0.82 ± 0.38
Pickup Above 0.02 ± 0.09 0.91 ± 0.07 0.90 ± 0.08 0.99 ± 0.09
Pickup Dist 0.10 ± 0.24 0.86 ± 0.21 0.90 ± 0.07 1.05 ± 0.09
Pickup Loc 0.08 ± 0.23 0.91 ± 0.04 0.86 ± 0.16 0.94 ± 0.19
Put Next S7N4 0.00 ± 0.03 0.96 ± 0.01 0.85 ± 0.22 0.88 ± 0.23
Put Next Local 0.00 ± 0.05 0.92 ± 0.03 0.63 ± 0.36 0.69 ± 0.40
Synth 0.11 ± 0.26 0.93 ± 0.06 0.77 ± 0.33 0.80 ± 0.40
Synth Loc 0.13 ± 0.29 0.94 ± 0.06 0.79 ± 0.33 0.81 ± 0.40
Synth Seq 0.07 ± 0.20 0.95 ± 0.04 0.52 ± 0.44 0.52 ± 0.50
Unblock Pickup 0.08 ± 0.22 0.91 ± 0.08 0.74 ± 0.32 0.79 ± 0.39
Unlock 0.03 ± 0.15 0.87 ± 0.10 0.52 ± 0.42 0.58 ± 0.50
@@
Unlock Pickup 0.00 ± 0.00 0.75 ± 0.04 0.76 ± 0.04 1.01 ± 0.05
Unlock To Unlock 0.00 ± 0.00 0.96 ± 0.00 0.80 ± 0.35 0.83 ± 0.37
Figure 8: Expert normalized episodic return for the JAT agent on the BabyAI benchmark.
@@
BinPicking 1.9 ± 0.4 425.6 ± 101.9 200.0 ± 222.1 0.47 ± 0.52
Box Close 76.4 ± 17.9 512.5 ± 107.8 462.6 ± 172.0 0.89 ± 0.39
Button Press 31.7 ± 5.2 643.1 ± 12.8 560.0 ± 182.6 0.86 ± 0.30
Button Press Topdown 29.0 ± 10.4 490.2 ± 27.2 266.2 ± 77.2 0.51 ± 0.17
Button Press Topdown Wall 29.0 ± 10.5 497.2 ± 31.4 275.8 ± 88.6 0.53 ± 0.19
Button Press Wall 9.0 ± 4.0 675.4 ± 15.0 638.3 ± 123.0 0.94 ± 0.18
Coffee Button 31.7 ± 6.4 731.1 ± 29.3 298.0 ± 285.6 0.38 ± 0.41
Coffee Pull 4.1 ± 0.4 259.9 ± 88.5 41.0 ± 69.8 0.14 ± 0.27
Coffee Push 4.2 ± 0.8 496.8 ± 118.2 153.1 ± 218.9 0.30 ± 0.44
Dial Turn 29.6 ± 16.7 793.6 ± 80.1 758.4 ± 120.4 0.95 ± 0.16
Disassemble 40.3 ± 7.5 42.8 ± 6.3 40.7 ± 9.9 0.17 ± 3.91
Door Close 5.3 ± 1.3 529.7 ± 27.2 524.3 ± 33.2 0.99 ± 0.06
Door Lock 112.3 ± 28.6 811.5 ± 34.1 696.3 ± 198.6 0.84 ± 0.28
Door Open 56.4 ± 11.2 581.9 ± 19.7 577.5 ± 53.7 0.99 ± 0.10
Door Unlock 94.2 ± 15.6 802.9 ± 17.1 768.3 ± 91.8 0.95 ± 0.13
Drawer Close 116.7 ± 253.1 867.9 ± 4.5 596.7 ± 223.4 0.64 ± 0.30
Drawer Open 126.8 ± 25.2 493.0 ± 2.5 485.9 ± 36.8 0.98 ± 0.10
Faucet Close 253.1 ± 22.9 753.9 ± 13.4 367.8 ± 91.1 0.23 ± 0.18
Faucet Open 244.1 ± 23.3 705.8 ± 7.1 566.1 ± 169.9 0.70 ± 0.37
Hammer 95.3 ± 9.0 693.2 ± 34.6 667.7 ± 89.4 0.96 ± 0.15
Hand Insert 2.8 ± 3.5 740.5 ± 36.7 688.1 ± 187.7 0.93 ± 0.25
Handle Press 80.4 ± 110.2 855.9 ± 72.7 735.0 ± 252.0 0.84 ± 0.32
Handle Press Side 57.0 ± 39.5 861.1 ± 20.0 64.5 ± 73.7 0.01 ± 0.09
Handle Pull 10.3 ± 13.5 669.4 ± 24.8 556.6 ± 161.7 0.83 ± 0.25
Handle Pull Side 2.1 ± 2.8 384.7 ± 102.9 195.1 ± 187.2 0.50 ± 0.49
Lever Pull 60.3 ± 15.8 612.0 ± 38.9 280.9 ± 234.8 0.40 ± 0.43
Peg Insert Side 1.7 ± 0.4 315.2 ± 140.1 254.3 ± 158.4 0.81 ± 0.51
Peg Unplug Side 4.7 ± 2.8 456.1 ± 81.7 80.6 ± 145.5 0.17 ± 0.32
@@
Pick Place 1.6 ± 1.0 419.1 ± 98.2 135.8 ± 200.1 0.32 ± 0.48
Pick Place Wall 0.0 ± 0.0 450.6 ± 64.1 43.7 ± 129.7 0.10 ± 0.29
Plate Slide 74.6 ± 13.8 527.0 ± 155.3 481.5 ± 190.2 0.90 ± 0.42
Plate Slide Back 33.5 ± 11.2 718.2 ± 87.4 196.9 ± 1.7 0.24 ± 0.00
Plate Slide Back Side 34.3 ± 11.5 729.6 ± 69.1 703.7 ± 117.3 0.96 ± 0.17
Plate Slide Side 22.6 ± 17.4 662.8 ± 102.8 122.6 ± 24.6 0.16 ± 0.04
Push 5.5 ± 2.4 750.6 ± 44.0 702.4 ± 157.6 0.94 ± 0.21
Push Back 1.2 ± 0.2 85.0 ± 107.1 82.2 ± 108.0 0.97 ± 1.29
Push Wall 6.1 ± 3.2 748.9 ± 10.6 158.8 ± 224.6 0.21 ± 0.30
Reach 149.7 ± 44.7 681.4 ± 133.7 332.2 ± 171.5 0.34 ± 0.32
Reach Wall 143.3 ± 36.6 746.1 ± 104.2 631.6 ± 224.0 0.81 ± 0.37
Shelf Place 0.0 ± 0.0 241.3 ± 24.6 92.1 ± 112.0 0.38 ± 0.46
Soccer 5.7 ± 4.6 375.2 ± 140.2 291.6 ± 161.8 0.77 ± 0.44
Stick Pull 2.6 ± 1.4 523.6 ± 18.9 480.1 ± 119.3 0.92 ± 0.23
Stick Push 2.8 ± 1.0 627.9 ± 10.2 303.2 ± 298.8 0.48 ± 0.48
Sweep 11.2 ± 7.3 494.8 ± 43.3 16.7 ± 18.7 0.01 ± 0.04
Sweep Into 12.5 ± 10.7 799.2 ± 19.1 793.3 ± 47.5 0.99 ± 0.06
Window Close 57.5 ± 7.1 591.3 ± 38.6 414.3 ± 207.4 0.67 ± 0.39
Window Open 43.4 ± 2.1 590.8 ± 57.1 577.3 ± 63.9 0.98 ± 0.12
Figure 9: Expert normalized episodic return for the JAT agent on the Meta-World benchmark.
@@
Ant -59.9 ± 99.6 5846.4 ± 942.6 5110.5 ± 1720.8 0.88 ± 0.29
Inverted Double Pendulum 57.5 ± 17.5 9338.7 ± 352.6 8663.7 ± 1259.4 0.93 ± 0.14
Half Cheetah -285.0 ± 79.8 7437.8 ± 173.3 6595.9 ± 244.4 0.89 ± 0.03
Hopper 18.4 ± 17.1 1858.7 ± 534.1 1409.0 ± 385.6 0.76 ± 0.21
Humanoid 122.0 ± 35.3 6281.0 ± 1795.8 712.6 ± 120.6 0.10 ± 0.02
Inverted Pendulum 6.1 ± 3.5 475.4 ± 179.0 117.4 ± 22.0 0.24 ± 0.05
Pusher -149.7 ± 7.4 -25.2 ± 6.7 -25.0 ± 6.3 1.00 ± 0.05
Reacher -43.0 ± 3.9 -5.7 ± 2.5 -5.9 ± 2.4 0.99 ± 0.06
Humanoid Standup 33135.8 ± 2481.9 273574.2 ± 85253.3 116736.7 ± 22765.5 0.35 ± 0.09
Swimmer 0.8 ± 10.7 92.2 ± 4.4 94.0 ± 4.1 1.02 ± 0.04
Walker 2d 2.7 ± 6.1 4631.2 ± 1059.0 4381.3 ± 851.1 0.95 ± 0.18
Figure 10: Expert normalized episodic return for the JAT agent on the MuJoCo benchmark.
@@ Appendix C Image captioning additional examples
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption